### v1

@@ 3.2 A posteriori error estimates in H(curl)
(3.7a) $a(u,v)=({\rm curl}u,{\rm curl}v)+(u,v)$ $=(f,v),\quad v\in H({\rm curl},\Omega),$
(3.7b) $\langle A_{h}({\rm curl})u_{h},v\rangle=({\rm curl}u_{h},{\rm curl}v)+(u_{h},v)$ $=(f,v),\quad v\in V_{h}=V_{h}(\rm curl),$
@@
Let $I_{h}:V_{h}({\rm curl})\hookrightarrow V_{h/2}({\rm curl})$ be the inclusion. It has been shown in [3] that
@@
(3.8) $\|u-u_{h}\|_{H({\rm curl},\Omega)}\eqsim\langle r,S_{a}({\rm curl})r\rangle^{\frac{1}{2}}=\Big(\sum_{k=1}^{N_{h/2}}\|\eta_{k}\|_{H({\rm curl},\Omega_{h/2,k})}\Big)^{\frac{1}{2}}.$
@@
(3.9) $\|u-u_{h}\|_{H({\rm curl},\Omega)}\eqsim\|S_{m}({\rm curl})r\|_{H({\rm curl},\Omega)}.$
@@
Figure 5: Convergence of error estimators for $H({\rm curl})$ problems in 2D (left) and 3D (right).
Figure 6: Convergence of a posteriori error estimates for controlling the $L^{2}$ norm error in 2D (left) and 3D (right).
Figure 7: Adaptive meshes (approximately $2.6\times 10^{4}$ tetrahedra) on the 3D cavity region generated by error estimators controlling the $H({\rm curl})$ norm (left) and $L^{2}$ norm (right) errors.
The H(curl) problem (3.7) is discretized by the lowest order edge finite element on the L-shaped domain $\Omega=(-1,1)^{2}\backslash([0,1)\times[-1,0))$ as well as on the 3D cavity region $\Omega=(0,1)^{3}\backslash[0.25,0.75]^{3}$, see Figure 7. We compare the smoother-type a posteriori error estimates (3.8) and (3.9) with the residual-type error estimator (cf. [14, 51, 42])
@@
$\quad+\sum_{F\text{ is a face in }\mathcal{T}_{h}}h_{F}\|\llbracket(f-u_{h})\cdot n\rrbracket\|^{2}_{L^{2}(F)}+h_{F}\|\llbracket{\rm curl}u_{h}\wedge n\rrbracket\|^{2}_{L^{2}(F)}\Big)^{\frac{1}{2}}.$
Here ${\rm curl}u_{h}\wedge n|_{F}={\rm curl}u_{h}\cdot t$ in $\mathbb{R}^{2}$ with $t$ being the unit tangent vector along $F$ and ${\rm curl}u_{h}\wedge n|_{F}={\rm curl}\times n$ in $\mathbb{R}^{3}$ with $n$ being the unit normal to $F$. On the L-shaped domain $\Omega$, the exact solution is set to be $u=\nabla(r^{2/3}\sin(2\theta/3))$, where $r^{2/3}\sin(2\theta/3)$ is the singular solution used in section 3.1. On the 3D test region, we set $f=(\sqrt{x^{2}-2y^{2}+3z^{2}+3}\exp(-xyz),0,0)$ in (3.7) and use $\|u_{h}-u_{h/2}\|_{H(\rm curl,\Omega)}$ as the reference FEM error as the exact solution $u$ is unknown. It is shown in Figure 5 that all estimators lead to optimally convergent FEM solutions, while smoother-type a posteriori error estimates are much closer to the exact error than the residual error estimators.
We are also interested in a posteriori error control of the $L^{2}$ norm error $\|u-u_{h}\|_{L^{2}(\Omega)}$ of (3.7b) based on Nédélec’s edge finite elements of the second kind. In this scenario, $\|u-u_{h}\|_{L^{2}(\Omega)}$ converges at a rate one order higher than $\|{\rm curl}(u-u_{h})\|_{L^{2}(\Omega)}$. Currently, we are not aware of any residual-type a posteriori error estimate for $\|u-u_{h}\|_{L^{2}(\Omega)}$. Motivated by the $L^{2}$ norm error estimators in section 3.1, we employ $\|S_{a}({\rm curl})r\|_{L^{2}(\Omega)}$ to guide the local mesh refinement of the lowest order Nédélec’s edge element of the second kind. As shown in Figure 4, the AFEM driven by $\|S_{a}({\rm curl})r\|_{L^{2}(\Omega)}$ recovers optimal convergence rate in the $L^{2}$ norm.
@@ 4 Block Diagonal Estimators for Symmetric Systems
(4.1a) $a(\sigma,\tau)+b(\tau,u)$ $=g(\tau),\quad\tau\in\Sigma,$
(4.1b) $b(\sigma,v)$ $=f(v),\quad v\in U.$
Here $a:\Sigma\times\Sigma\rightarrow\mathbb{R}$, $b:\Sigma\times U\rightarrow\mathbb{R}$ are bounded bilinear forms, $a$ is symmetric and $f\in\Sigma^{*}$, $g\in U^{*}$. Let $\Sigma_{h}\times U_{h}\subset\Sigma\times U$ be an inf-sup stable finite element space pair equipped with the inner products $\langle A_{\Sigma_{h}}\bullet,\bullet\rangle$ and $\langle A_{U_{h}}\bullet,\bullet\rangle$, respectively, where $A_{\Sigma_{h}}:\Sigma_{h}\rightarrow\Sigma_{h}^{*}$ and $A_{U_{h}}:U_{h}\rightarrow U_{h}^{*}$ are SPD. A mixed FEM for (4.1) seeks $\sigma_{h}\in\Sigma_{h}$ and $u_{h}\in U_{h}$ satisfying
(4.2a) $a(\sigma_{h},\tau)+b(\tau,u_{h})$ $=g(\tau),\quad\tau\in\Sigma_{h},$
(4.2b) $b(\sigma_{h},v)$ $=f(v),\quad v\in U_{h}.$
@@
(4.3) $\|(\sigma-\sigma_{h/2},u-u_{h/2})\|_{\Sigma\times U}\leq\gamma_{h/2}\|(\sigma-\sigma_{h},u-u_{h})\|_{\Sigma\times U}$
@@
Assume that $\Sigma_{h}\subset\Sigma_{h/2}$ and $U_{h}\subset U_{h/2}$ with $I_{h}:\Sigma_{h}\rightarrow\Sigma_{h/2}$ and $J_{h}:U_{h}\rightarrow U_{h/2}$ being inclusions. Let $B_{\Sigma_{h/2}}=I_{h}A_{\Sigma_{h/2}}^{-1}I_{h}^{*}+S_{\Sigma_{h/2}}:\Sigma_{h/2}^{*}\rightarrow\Sigma_{h/2}$ and $B_{U}=J_{h}A_{U_{h/2}}^{-1}J_{h}^{*}+S_{U_{h/2}}:U_{h/2}^{*}\rightarrow U_{h/2}$. Assume that $B_{\Sigma_{h/2}}$, $B_{U_{h/2}}$ are preconditioners for $A_{\Sigma_{h/2}}$, $A_{U_{h/2}}$, respectively, i.e.,
@@
where $c_{2},...,c_{5}>0$ are uniform constants. Then it holds that
$\frac{\min(c_{2},c_{4})}{\alpha_{h/2}(1+\gamma_{h/2})}\mathcal{E}_{h}(\sigma_{h},u_{h})\leq\|(\sigma-\sigma,u-u_{h})\|_{\Sigma\times U}\leq\frac{\max(c_{3},c_{5})}{\beta_{h/2}(1-\gamma_{h/2})}\mathcal{E}_{h}(\sigma_{h},u_{h}).$
@@
Proof 4.2.
By definition of the inf-sup condition and $I_{h}^{*}r_{\sigma}=0$, $J_{h}^{*}r_{u}=0$, we have
@@
(4.4) $\|(\sigma_{h/2}-\sigma_{h},u_{h/2}-u_{h})\|_{\Sigma\times U}\leq\frac{\max(c_{3},c_{5})}{\beta_{h}}\sqrt{\langle r_{\sigma},S_{\Sigma_{h/2}}r_{\sigma}\rangle+\langle r_{u},S_{U_{h/2}}r_{u}\rangle}.$
@@
$\|(\sigma-\sigma_{h},u-u_{h})\|_{\Sigma\times U}\leq\frac{\max(c_{3},c_{5})}{\beta_{h}(1-\gamma_{h/2})}\sqrt{\langle r_{\sigma},S_{\Sigma_{h/2}}r_{\sigma}\rangle+\langle r_{u},S_{U_{h/2}}r_{u}\rangle}.$
The lower bound is proved in a similar way.
4.1 Poisson Euqation in Mixed Form
@@
In [3], it has been shown that
@@
is a uniform preconditioner for $A_{h/2}({\rm div})$. Therefore, we have verified the assumptions in Theorem 4.1 and obtain
(4.6) $\|\sigma-\sigma_{h}\|^{2}_{H({\rm div},\Omega)}+\|u-u_{h}\|^{2}_{L^{2}(\Omega)}\eqsim\langle r_{\sigma},S_{a}({\rm div})r_{\sigma}\rangle+\langle r_{u},A_{U_{h/2}}^{-1}r_{u}\rangle.$
Figure 8: Convergence of a posteriori error estimates for the mixed FEM (4.5).
@@
Figure 5 demonstrates that the smoother-type estimator (4.6) and is more accurate than $\mathcal{E}_{\rm res,mix}$ and yields an optimally convergent AFEM.
@@ 4.2 Biot’s Consolidation Model
(4.7a) $-{\rm div}(\varepsilon(u)+\lambda({\rm div}u)I-\alpha p_{F}I)$ $=f\quad\text{ in }\Omega,$
(4.7b) $-\alpha^{2}\lambda^{-1}p_{F}-\alpha{\rm div}u+{\rm div}(\kappa\nabla p_{F})$ $=g\quad\text{ in }\Omega,$
@@
(4.8a) $(\varepsilon(u_{h}),\varepsilon(v_{h}))-({\rm div}v_{h},p_{T,h})$ $=(f,v_{h}),$
(4.8b) $-({\rm div}u_{h},q_{T,h})-\lambda^{-1}(p_{T,h},q_{T,h})+\alpha\lambda^{-1}(p_{F,h},q_{T,h})$ $=0,$
(4.8c) $\alpha\lambda^{-1}(p_{T,h},q_{F,h})-2\alpha^{2}\lambda^{-1}(p_{F,h},q_{F,h})-\kappa(\nabla p_{F,h},\nabla q_{F,h})$ $=(g,q_{F})$
@@
Let $H_{0}^{1}(\Omega)]^{d}$, $q_{T}\in L_{0}^{2}(\Omega)$ and $H_{0}^{1}(\Omega)$ be equipped with the norms
@@
Let $S_{\widetilde{V}}:\widetilde{V}^{*}\rightarrow\widetilde{V}$ be the pointwise Jacobi smoother, namely, the operator corresponding to the inverse diagonal of the matrix for $A_{\widetilde{V}}$ with $\widetilde{V}=V_{h/2}$ or $Q_{F,h/2}$. Additive two-level preconditioners using the smoother $S_{\widetilde{V}}$ suffices to uniformly precondition the second order elliptic operator $A_{\widetilde{V}}$. Although $A_{Q_{T,h/2}}$ corresponds to a mass matrix, the global average in $A_{Q_{T,h/2}}$ makes the matrix for $A_{Q_{T,h/2}}$ dense. In section 5 of [38], $A_{Q_{T,{h/2}}}$ is preconditioned by $S_{Q_{T,h/2}}:Q_{T,h/2}^{*}\rightarrow Q_{T,h/2}$ corresponding to a matrix of the form $(\mathbb{I}+\mathbb{V})\mathbb{D}(\mathbb{I}+\mathbb{V}^{\top})$, where $\mathbb{I}$ is the identity, $\mathbb{D}$ is diagonal and $\mathbb{V}$ is of rank one.
@@
Figure 9: Uniform convergence of error estimators for (4.7) using a constructed solution with effectivity ratio estimator/error shown in the parenthesis.
Figure 10: Convergence of the AFEM driven by a smoother-type error estimator for (4.7) on a L-shaped domain with $f=(1,1)$, $g=1$ and ratio estimator/error shown in the parenthesis.
@@
as the reference error. The corresponding numerical results are presented in Figure 10. The smoother-type error estimators are able to accurately control the FEM error under locally refined meshes for singular solutions.

### v2

@@ 3.2 A Posteriori Error Estimates in H(curl)
$a(u,v)=({\rm curl}u,{\rm curl}v)+(u,v)$ $=(f,v),\quad v\in H({\rm curl},\Omega),$ (3.8a)
$\langle A_{h}({\rm curl})u_{h},v\rangle=({\rm curl}u_{h},{\rm curl}v)+(u_{h},v)$ $=(f,v),\quad v\in V_{h}=V_{h}(\rm curl),$ (3.8b)
@@
Let $I_{h}:V_{h}({\rm curl})\hookrightarrow V_{h/2}({\rm curl})$ be the inclusion. It has been shown in [4] that $V_{h/2}({\rm curl})=V_{h}({\rm curl})+\sum_{k=1}^{N_{h/2}}V_{k}({\rm curl})$ is a stable subspace decomposition and
@@
$\|u-u_{h}\|_{H({\rm curl},\Omega)}\eqsim\langle r,S_{a}({\rm curl})r\rangle^{\frac{1}{2}}=\Big(\sum_{k=1}^{N_{h/2}}\|\eta_{k}\|_{H({\rm curl},\Omega_{h/2,k})}\Big)^{\frac{1}{2}}.$ (3.9)
@@
$\|u-u_{h}\|_{H({\rm curl},\Omega)}\eqsim\|S_{m}({\rm curl})r\|_{H({\rm curl},\Omega)}.$ (3.10)
@@
Figure 3.5: Convergence of error estimators for $H({\rm curl})$ problems in 2D (left) and 3D (right).
Figure 3.6: Convergence of a posteriori error estimates for controlling the $L^{2}$ norm error in 2D (left) and 3D (right).
Figure 3.7: Adaptive meshes (approximately $2.6\times 10^{4}$ tetrahedra) on the 3D cavity region generated by error estimators controlling the $H({\rm curl})$ norm (left) and $L^{2}$ norm (right) errors.
The H(curl) problem (3.8) is discretized by the lowest order edge finite element on the L-shaped domain $\Omega=(-1,1)^{2}\backslash([0,1)\times[-1,0))$ as well as on the 3D cavity region $\Omega=(0,1)^{3}\backslash[0.25,0.75]^{3}$, see Figure 3.7. We compare the smoother-type a posteriori error estimates (3.9) and (3.10) with the residual-type error estimator (cf. [16, 60, 52])
@@
$\quad+\sum_{F\text{ is a face in }\mathcal{T}_{h}}h_{F}\|\llbracket(f-u_{h})\cdot n\rrbracket\|^{2}_{L^{2}(F)}+h_{F}\|\llbracket n\wedge{\rm curl}u_{h}\rrbracket\|^{2}_{L^{2}(F)}\Big)^{\frac{1}{2}}.$
Here $n\wedge{\rm curl}u_{h}|_{F}=t\cdot{\rm curl}u_{h}$ in $\mathbb{R}^{2}$ with $t$ being the unit tangent vector along $F$ and $n\wedge{\rm curl}u_{h}|_{F}=n\times{\rm curl}u_{h}$ in $\mathbb{R}^{3}$ with $n$ being the unit normal to $F$. On the L-shaped domain $\Omega$, the exact solution is set to be $u=\nabla(r^{2/3}\sin(2\theta/3))$, where $r^{2/3}\sin(2\theta/3)$ is the singular solution used in Section 3.1. On the 3D test region, we set $f=(\sqrt{x^{2}-2y^{2}+3z^{2}+3}\exp(-xyz),0,0)$ in (3.8) and use $\|u_{h}-u_{h/2}\|_{H(\rm curl,\Omega)}$ as the reference FEM error as the exact solution $u$ is unknown. It is shown in Figure 3.5 that all estimators lead to optimally convergent FEM solutions, while smoother-type a posteriori error estimates are much closer to the exact error than the residual error estimators.
We are also interested in a posteriori error control of the $L^{2}$ norm error $\|u-u_{h}\|_{L^{2}(\Omega)}$ of (3.8b) based on Nédélec’s edge finite elements of the second kind. In this scenario, $\|u-u_{h}\|_{L^{2}(\Omega)}$ converges at a rate one order higher than $\|{\rm curl}(u-u_{h})\|_{L^{2}(\Omega)}$. Currently, we are not aware of any residual-type a posteriori error estimate for $\|u-u_{h}\|_{L^{2}(\Omega)}$. Motivated by the $L^{2}$ norm error estimators in Section 3.1, we employ $\|S_{a}({\rm curl})r\|_{L^{2}(\Omega)}$ to guide the local mesh refinement of the lowest order Nédélec’s edge element of the second kind. As shown in Figure 3.4, the AFEM driven by $\|S_{a}({\rm curl})r\|_{L^{2}(\Omega)}$ recovers optimal convergence rate in the $L^{2}$ norm.
@@ 4 Block Diagonal Estimators for Symmetric Systems
$a(\sigma,\tau)+b(\tau,u)$ $=g(\tau),\quad\tau\in\Sigma,$ (4.1a)
$b(\sigma,v)$ $=f(v),\quad v\in U.$ (4.1b)
Here $a:\Sigma\times\Sigma\rightarrow\mathbb{R}$, $b:\Sigma\times U\rightarrow\mathbb{R}$ are bounded bilinear forms, $a$ is symmetric and $f\in U^{*}$, $g\in\Sigma^{*}$. Let $\Sigma_{h}\times U_{h}\subset\Sigma\times U$ be an inf-sup stable finite element space pair equipped with the inner products $\langle A_{\Sigma_{h}}\bullet,\bullet\rangle$ and $\langle A_{U_{h}}\bullet,\bullet\rangle$, respectively, where $A_{\Sigma_{h}}:\Sigma_{h}\rightarrow\Sigma_{h}^{*}$ and $A_{U_{h}}:U_{h}\rightarrow U_{h}^{*}$ are SPD. A mixed FEM for (4.1) seeks $\sigma_{h}\in\Sigma_{h}$ and $u_{h}\in U_{h}$ satisfying
$a(\sigma_{h},\tau)+b(\tau,u_{h})$ $=g(\tau),\quad\tau\in\Sigma_{h},$ (4.2a)
$b(\sigma_{h},v)$ $=f(v),\quad v\in U_{h}.$ (4.2b)
@@
$\|(\sigma-\sigma_{h/2},u-u_{h/2})\|_{\Sigma\times U}\leq\gamma_{h/2}\|(\sigma-\sigma_{h},u-u_{h})\|_{\Sigma\times U}$ (4.3)
@@
Assume that $\Sigma_{h}\subset\Sigma_{h/2}$ and $U_{h}\subset U_{h/2}$ with $I_{h}:\Sigma_{h}\rightarrow\Sigma_{h/2}$ and $J_{h}:U_{h}\rightarrow U_{h/2}$ being inclusions. Let $B_{\Sigma_{h/2}}=I_{h}A_{\Sigma_{h/2}}^{-1}I_{h}^{*}+S_{\Sigma_{h/2}}:\Sigma_{h/2}^{*}\rightarrow\Sigma_{h/2}$ and $B_{U_{h/2}}=J_{h}A_{U_{h/2}}^{-1}J_{h}^{*}+S_{U_{h/2}}:U_{h/2}^{*}\rightarrow U_{h/2}$. Assume that $B_{\Sigma_{h/2}}$, $B_{U_{h/2}}$ are preconditioners for $A_{\Sigma_{h/2}}$, $A_{U_{h/2}}$, respectively, i.e.,
@@
where $c_{2},...,c_{5}>0$ are uniform constants. Then for (4.2) it holds that
$\frac{\sqrt{\min(c_{2},c_{4})}}{\alpha_{h/2}(1+\gamma_{h/2})}\mathcal{E}_{h}(\sigma_{h},u_{h})\leq\|(\sigma-{\sigma_{h}},u-u_{h})\|_{\Sigma\times U}\leq\frac{\sqrt{\max(c_{3},c_{5})}}{\beta_{h/2}(1-\gamma_{h/2})}\mathcal{E}_{h}(\sigma_{h},u_{h}).$
@@
Proof.
Using $\Sigma_{h}\subset\Sigma_{h/2}$, $U_{h}\subset U_{h/2}$, we have $I_{h}^{*}r_{\sigma}=0$, $J_{h}^{*}r_{u}=0$ and thus $\langle r_{\sigma},B_{\Sigma_{h/2}}r_{\sigma}\rangle=\langle r_{\sigma},S_{\Sigma_{h/2}}r_{\sigma}\rangle$ and $\langle r_{u},B_{U_{h/2}}r_{u}\rangle=\langle r_{u},S_{U_{h/2}}r_{u}\rangle$. By definition of the inf-sup condition and spectral equivalence assumption, we have
@@
$\|(\sigma_{h/2}-\sigma_{h},u_{h/2}-u_{h})\|_{\Sigma\times U}\leq\frac{\sqrt{\max(c_{3},c_{5})}}{\beta_{h}}\sqrt{\langle r_{\sigma},S_{\Sigma_{h/2}}r_{\sigma}\rangle+\langle r_{u},S_{U_{h/2}}r_{u}\rangle}.$ (4.4)
@@
$\|(\sigma-\sigma_{h},u-u_{h})\|_{\Sigma\times U}\leq\frac{\sqrt{\max(c_{3},c_{5})}}{\beta_{h}(1-\gamma_{h/2})}\sqrt{\langle r_{\sigma},S_{\Sigma_{h/2}}r_{\sigma}\rangle+\langle r_{u},S_{U_{h/2}}r_{u}\rangle}.$
The lower bound is proved in a similar way. ∎
4.1 Poisson in Mixed Form
@@
In [4], it has been shown that $V_{h/2}({\rm div})=V_{h}({\rm div})+\sum_{k=1}^{N_{h/2}}V_{k}({\rm div})$ is a stable subspace decomposition and
@@
is a uniform preconditioner for the discrete $H({\rm div})$ operator $A_{h/2}({\rm div})$. Therefore, for the mixed FEM (4.5), we have verified the assumptions in Theorem 4.1 and obtain
$\|\sigma-\sigma_{h}\|^{2}_{H({\rm div},\Omega)}+\|u-u_{h}\|^{2}_{L^{2}(\Omega)}\eqsim\langle r_{\sigma},S_{a}({\rm div})r_{\sigma}\rangle+\langle r_{u},A_{U_{h/2}}^{-1}r_{u}\rangle.$ (4.6)
Figure 4.1: Convergence of a posteriori error estimates for the mixed FEM (4.5).
@@
Figure 4.1 demonstrates that the smoother-type estimator (4.6) is more accurate than $\mathcal{E}_{\rm res,mix}$ and yields an optimally convergent AFEM.
@@ 4.2 Biot’s Consolidation Model
$-{\rm div}(\varepsilon(u)+\lambda({\rm div}u)I-\alpha p_{F}I)$ $=f\quad\text{ in }\Omega,$ (4.7a)
$-\alpha^{2}\lambda^{-1}p_{F}-\alpha{\rm div}u+{\rm div}(\kappa\nabla p_{F})$ $=g\quad\text{ in }\Omega,$ (4.7b)
@@
$(\varepsilon(u_{h}),\varepsilon(v_{h}))-({\rm div}v_{h},p_{T,h})$ $=(f,v_{h}),$ (4.8a)
$-({\rm div}u_{h},q_{T,h})-\lambda^{-1}(p_{T,h},q_{T,h})+\alpha\lambda^{-1}(p_{F,h},q_{T,h})$ $=0,$ (4.8b)
$\alpha\lambda^{-1}(p_{T,h},q_{F,h})-2\alpha^{2}\lambda^{-1}(p_{F,h},q_{F,h})-\kappa(\nabla p_{F,h},\nabla q_{F,h})$ $=(g,{q_{F,h}})$ (4.8c)
@@
Let $H_{0}^{1}(\Omega)]^{d}$, $L_{0}^{2}(\Omega)$ and $H_{0}^{1}(\Omega)$ be equipped with the norms
@@
Let $S_{\widetilde{V}}:\widetilde{V}^{*}\rightarrow\widetilde{V}$ be the pointwise Jacobi smoother, namely, the operator corresponding to the inverse diagonal of the matrix for $A_{\widetilde{V}}$ with $\widetilde{V}=V_{h/2}$ or $Q_{F,h/2}$. Additive two-level preconditioners using the smoother $S_{\widetilde{V}}$ suffice to uniformly precondition the second order elliptic operator $A_{\widetilde{V}}$. Although $A_{Q_{T,h/2}}$ corresponds to a mass matrix, the global average in $A_{Q_{T,h/2}}$ makes the matrix for $A_{Q_{T,h/2}}$ dense. In Section 5 of [47], $A_{Q_{T,{h/2}}}$ is preconditioned by $S_{Q_{T,h/2}}:Q_{T,h/2}^{*}\rightarrow Q_{T,h/2}$ corresponding to a matrix of the form $(\mathbb{I}+\mathbb{V})\mathbb{D}(\mathbb{I}+\mathbb{V}^{\top})$, where $\mathbb{I}$ is the identity, $\mathbb{D}$ is diagonal and $\mathbb{V}$ is of rank one.
@@
Figure 4.2: Uniform convergence of error estimators for (4.7) using a constructed solution with effectivity ratio estimator/error shown in the parenthesis.
Figure 4.3: Convergence of the AFEM driven by a smoother-type error estimator for (4.7) on an L-shaped domain with $f=(1,1)$, $g=1$ and ratio estimator/error shown in the parenthesis.
@@
as the reference error. The corresponding numerical results are presented in Figure 4.3. The smoother-type error estimators are able to accurately control the FEM error under locally refined meshes for singular solutions.